# Human4DiT: Free-view Human Video Generation with 4D Diffusion Transformer

### v1

@@ 7.3 Ablation Study
To validate the effectiveness of introducing the fourth view dimension and the view transformer in our 4D transformer, we conducted ablation studies for multi-view video, 3D static video, and free-view video generation tasks. Specifically, we disabled the view transformer and relied solely on the temporal transformer to generate the corresponding videos. Quantitative comparisons are presented in Tab. 3, 4, 5. It can be observed that the introduction of the view transformer leads to improvements in view-related generation tasks, demonstrating the efficacy of the view transformer.
8 Discussion
@@
Conclusion. We have presented a novel method for human video generation, which takes as input only a single image and produces spatio-temporally coherent video of dynamic human motions under free viewpoints. Our approach employs an efficient 4D transformer architecture to model the correlations across multiple domains, including view, time and poses. Combining with UNets for accurate condition injection, our model can be trained on a multi-dimensional dataset spanning images, videos, multi-view data, and 4D scans. After training, our method can synthesize realistic, coherent human motion videos from free viewpoints, and we believe our contributions will inspire future work towards 4D content generation.
@@
Limitations. Instead of generating an explicit 4D models, our method directly synthesizes 2D videos from the given viewpoints, and the 4D scene structure are implicitly encoded through the attention mechanism. The absence of an explicit 4D representation results in some subtle artifacts when rendering free-view videos. Furthermore, our current implementation cannot generate tiny structures coherently, such as fingers and accessories. Future works may incorporate body part-aware generator to resolve this limitation.

### v2

@@ 7.3.2. View Transformer
To validate the effectiveness of introducing the fourth view dimension and the view transformer in our 4D transformer, we conduct ablation studies for multi-view video, 3D static video, and 360-degree video generation tasks. Specifically, we disable the view transformer and rely solely on the temporal transformer to generate the corresponding videos. Quantitative results of ”Ours (image+temporal)” are presented in Tab. 3.It can be observed that the introduction of the view transformer leads to improvements in view-related generation tasks, demonstrating the efficacy of the view transformer.
7.3.3. Efficient Spatial-temporal Sampling
To validate the efficacy of our spatial-temporal sampling, we conducted a qualitative comparison, as illustrated in Fig 5. When only temporal window sampling is employed, the generated 360-degree video exhibits noticeable misalignments (The tie on the person’s chest has become a flat, texture surface). In contrast, our spatial-temporal sampling method mitigates these inconsistencies, achieving global coherence throughout the generated sequence. The results demonstrate that our method effectively addresses the challenges of long-range dependencies in 360-degree human video generation, contributing to enhanced visual quality and coherence.
7.4. Failure Cases
As illustrated in Fig. 7, our work primarily addresses full-body, 360-degree human video generation. However, the face and hands of human, typically occupying less than 100x100 pixels within our 768x768 generated video frames, are leading to artifacts due to their relatively small scale. Furthermore, inaccuracies in SMPL estimation also affect the overall generation quality. We believe that future research incorporating specialized high-resolution modules for facial and hand generation could potentially address these issues.
Additionally, our current approach demonstrates limitations in effectively modeling background dynamics. This problem arises from the nature of our dataset, which predominantly features static backgrounds. Moreover, our 3D/4D datasets are rendered against a uniform white background. Consequently, the model has not learned to accurately represent background transformations across varying viewpoints.
8. Discussion
@@
Conclusion. We have presented a novel method for human video generation, which takes as input only a single image and produces spatio-temporally coherent video of dynamic human motions under 360-degree viewpoints. Our approach employs an efficient 4D transformer architecture to model the correlations across multiple domains, including view, time and poses. Combining with UNets for accurate condition injection, our model can be trained on a multi-dimensional dataset spanning images, videos, multi-view data, and 4D scans. After training, our method can synthesize 360-degree realistic, coherent human motion videos, and we believe our contributions will inspire future work towards 4D content generation.
@@
Limitations and Future Work. Instead of generating an explicit 4D models, our method directly synthesizes 2D videos from the given viewpoints, and the 4D scene structure are implicitly encoded through the attention mechanism. The absence of an explicit 4D representation results in some artifacts when rendering 360-degree videos. Furthermore, our current implementation cannot generate tiny structures coherently as mentioned in Sec. 7.4. Future works may incorporate a body part-aware generator to resolve this limitation. Exploring other condition injection techniques is also significant to enhance controllability. Potential directions include improved camera representations, Plücker embeddings, and direct utilization of 3D SMPL as tokens. We also believe collecting a more diverse dataset is significant to fully leverage the capabilities of DiT, and address video generation challenges in out-of-distribution (OOD) scenarios, such as anime or stylized content.
9. Ethics Statement
While our method for human video generation could benefit a lot of applications for entertainment, education, and accessibility, we recognize the risks of misuse for creating deepfakes, privacy violations, and perpetuating societal biases. We are committed to responsible development and promoting transparency in our methods. By discussing these ethical considerations, we aim to foster the positive applications of this technology while minimizing potential harm.
Acknowledgement The work is supported by the National Science Foundation of China under Grant Number 62125107.